Earned and Unearned Income (LO. 2)

Duc has been employed by Longbow Corporation for 25 years. During that time, he bought an annuity at a cost of $50 per month ($15,000 total cost). The annuity will pay him $200 per month after he reaches age 65. When Duc dies, his wife, Annika, will continue to receive the annuity until her death. Duc turns 65 in April 2019 and receives 8 payments on the contract. Annika is age 60 when the annuity payments begin. Refer to the Annuity payment tables to answer the following questions.

Round intermediate calculations to the nearest dollar.

a. Duc's gross income from the contract in the current year should be $._1216_

b. Assume that Duc dies on April 2, 2029. Annika will include $_1824_ in gross income for the contract in 2029.

c. Assume the same facts as in part b and that Annika dies on August 4, 2036. The executor of Annika's estate deducts $______ and includes $______ in gross income for the contract in the year of her death.
Answer #1

Answer-a:

15,000/310 = 48
200 - 48 = 152
152 * 8 = 1,216

Answer-b:

Annika receives the payments after Duc dies. She still excludes the $48 from taxes
152 * 12 = 1,824

Answer-c:

Executor includes 1,216 in her gross income (152 * 8)
Has only been paid 208 out of 310 payments (so not all of the 15,000) have only excluded 9,984 from their contract (48 * 208)
Entitled to a deduction of 5,016 from her return (15,000 - 9,984)